A sinusoidal wave is described by 𝑦 = (0.48 π‘š) sin(0.555π‘₯ βˆ’ 36.0𝑑) π‘š, where x and y are in meters


and t is in seconds. Determine for this wave the (a) amplitude, (b) angular frequency, (c) angular wave


number, (d) wavelength, (e) wave speed, and (f) direction of motion.



1
Expert's answer
2021-12-08T10:00:23-0500

(a) "A=0.48\ m."

(b) "\omega=36\ \dfrac{rad}{s}."

(c) "k=0.555\ m^{-1}."

(d)

"k=\dfrac{2\pi}{\lambda},""\lambda=\dfrac{2\pi}{k}=\dfrac{2\pi}{0.555\ m^{-1}}=11.32\ m."


(e)

"\omega=2\pi f,""f=\dfrac{\omega}{2\pi}=\dfrac{36\ \dfrac{rad}{s}}{2\pi}=5.73\ Hz,""v=f\lambda=5.73\ Hz\times11.32\ m=64.8\ \dfrac{m}{s}."

(f) The wave moving in negative "x"-direction.